# The Coase Theorem and Ideal Exchanges

### v4

@@ 1. Introduction
An externality can be defined as an effect, often negative, that is imposed on a third-party through a private activity. An instance of this problem occurs when non-smokers and smokers are within each other’s vicinity; the non-smoker is adversely affected by the smoker’s consumption of cigarettes through no fault of their own. The economic and philosophical significance of the existence of externalities lies in the challenge it poses to the classical position that a free market economy is capable of allocating resources in a socially efficient manner; if the social costs of private consumption are not reflected in the market price of the product in question, then consumers would be incentivized to consume the good at a level that is beyond socially optimal. A comprehensive overview of the economic explanations surrounding externalities and their historical origins can be found in [1].
Prior to Ronald Coase’s work, The Problem of Social Cost, the conventional view—which Coase himself attributes to Arthur Pigou [4, p. 1]—held that externalities could be corrected through state intervention: namely, through taxes or subsidies that are designed to deter people from either over-consuming goods associated with negative externalities or under-consuming goods associated with positive externalities [11]. Coase reoriented the debate from one that relied solely on governmental regulation to one that recognized the potential efficiency of private negotiation in resolving externalities. In particular, Coase reasoned that externalities resulted from ill-defined property rights; provided that property rights are well-defined and each party could negotiate with the other without hindrance, resources would be distributed away from those who value them less and towards those who value them more irrespective of how the rights are initially distributed [4, p. 8, para. 2]. This proposition, although informal, was characterized as a theorem in George Stigler’s Theory of Price:
@@
Returning to the initial example, if the non-smoker has the right to clean air, the smoker would have to compensate them to smoke in shared spaces. Conversely, if the smoker has the right to smoke, the non-smoker might compensate the smoker to refrain from smoking near them. Hence, regardless of which party initially holds the right, the outcome of these negotiations will be efficient, as the final allocation of rights will reflect the true costs and benefits to both sides.
@@
Stigler’s recognition of the potential mathematical standing of what is now widely termed the Coase theorem invites us to examine the economic proposition precisely, though an exact formulation is not present in Coase’s original work. The prevailing interpretation of the theorem adopted by this paper can be stated as follows: if property rights are well-defined and there are no transaction costs, then rational agents would arrive at an optimal distribution of resources on their own accord. We find this interpretation in Stigler [12], Nutter [10], Demsetz [7], and Coase himself [4]. Yet, again, we are confronted with the difficulty of constructing a precise interpretation of informal notions such as the property of being “well-defined” or constituting a “transaction cost”. Additionally, we may interpret the notion of “arriving at an optimal outcome” as a declaration of the existence of some ultimate, optimal distribution such that it is a necessary outcome of every possible initial distribution; alternatively, we may consider that there are multiple ultimate distributions for each initial distribution that fall within the set of outcomes characterized as being optimal.
This paper endeavors to formalize interpretations of the Coase theorem as set-theoretic propositions. We offer the philosophical motivations for doing this in Section 2 and construct the formal aspects of the theorem thereafter. Where necessary, explanatory remarks are offered to justify the formalization. We show that certain interpretations of the theorem are false whereas one is true if we admit the notion of ideal exchanges. Lastly, we conclude by offering some philosophical remarks on these results in Section 6.
@@ 2. Preliminaries
Contemporary methods of illustrating the Coase theorem are either informal or rely on mathematical structures that are unnatural within the discipline of economics. We shall concern ourselves with the latter. One such illustration can be obtained by employing the standard techniques of an indifference curve analysis, which are readily available in any microeconomics textbook.
Consider the existence of two types of goods, $x$ and $y$, such that $Q_{x}$ denotes the quantity of the former and $Q_{y}$ denotes the quantity of the latter. It is standard to assume that each 2-tuple $(Q_{x},Q_{y})$ is present in $\mathbb{R}^{+}\times\mathbb{R}^{+}$ or, equivalently, +$\mathbb{R}^{2}_{+}$. [^1] We include 00 in $\mathbb{R}^{+}$. We proceed by declaring the existence of a utility function for some agent $A$ such that $\mathcal{U}_{A}:\mathbb{R}^{2}_{+}\to\mathbb{R}^{+}$ to model the notion that it is possible to map pairs of quantities of goods to some level of utility. [^2] It would seem that happiness is a crude synonym for utility, since the latter is often portrayed as a quantifiable notion. An agent is indifferent between two outcomes iff they render the same level of utility. In other words, an indifference curve is a set of all the 2-tuples $(Q_{x},Q_{y})\in\mathbb{R}^{2}_{+}$ such that $\mathcal{U}_{A}(Q_{x},Q_{y})=c$, where $c$ is a constant in $\mathbb{R}^{+}$. Furthermore, it is standard for an economist to assume that each indifference curve carries important characteristics. Intuitively speaking, it would be peculiar for an indifference curve to contain $(Q_{x},Q_{y})$ and $(Q_{x}^{\prime},Q_{y}^{\prime})$ if $Q_{x}<Q_{x}^{\prime}$ and $Q_{y}<Q_{y}^{\prime}$, since that would imply that the utility of the agent remains constant despite an increase in the presence of both goods. It follows, therefore, that all indifference curves must be monotonic and decreasing. A more refined assumption would be that indifference curves should model the empirical observation of diminishing marginal utility, that is, every additional unit of utility gained from each additional unit of the good decreases, but never goes below zero. [^3] If it did, then that would imply that an additional unit of the good resulted in a loss of utility, which is not a standard assumption in microeconomic theory. Thus, indifference curves are often portrayed as being convex to the agent’s origin.
@@
Every Pareto optimal distribution occurs when the indifference curves for both agents are tangent to each other. We can see this by considering the intersection at $\omega_{A}$ or, equivalently, $\omega_{B}$. Since $A$ has an excess of good $x$ and $B$ has an excess of good $y$, it would be possible for $A$ to give up just enough of $x$ and receive just enough of $y$ so that she is indifferent between $\omega_{A}$ and $\Omega$, while $B$’s indifference curve progresses to a point where it is no longer possible to gain any further utility without damaging $A$’s material interests; the opposite is true at $\omega_{A}^{\prime}$. If we assume that exchanges can occur iff they are mutually beneficial, then the negotiated outcome is strictly inside the lens formed by the two indifference curves. Through continual negotiations, the size of the lens shrinks until it is no longer possible to make a mutually beneficial trade. [^4] This is a standard description of how negotiations occur. It is notably present in Buchanan and Tullock’s foundational work on public choice theory, The Calculus of Consent: Logical Foundations of Constitutional Democracy [2, p. 100]. Hence, the ultimate outcome, although different from $\Omega$, would still be Pareto optimal. A stronger variant of the theorem in narrow cases where a medium of exchange is involved was proven by Leonid Hurwicz, who demonstrated that the invariance of the final distribution of a particular resource necessarily depended on the assumption of quasi-linear preferences (i.e., $\mathcal{U}_{A}(Q_{x},m)=v(Q_{x})+m$, where $m$ is a “numeraire” good) [9].
@@
It is a striking fact that the preceding methods are standard in microeconomic theory yet are obviously mathematically unnatural for the purposes of modeling economic behavior. Francis Ysidro Edgeworth, after whom the Edgeworth box is named, introduced the concept of indifference curves in Mathematical Psychics (MP), a foundational work in contemporary microeconomic theory [8, p. 28]. The broad purpose of MP was to apply mathematical techniques to the moral sciences; in doing so, Edgeworth introduced utility measurements by assigning real numbers to represent individuals’ levels of satisfaction or happiness, facilitating the comparison and aggregation of utility across different people. Furthermore, MP readily and implicitly utilizes the concept of a continuum—suggesting infinitely divisible goods or utility levels—without being concerned with foundational aspects of real analysis. On a historical note, Dedekind’s Stetigkeit und irrationale Zahlen was published in 1872, merely 9 years before MP [6]; it is therefore likely that Edgeworth was unaware of these developments in the philosophical foundations of mathematics and their effects on the mathematical foundations of economics.
A mathematically bizarre example of such an effect is as follows; suppose that the utility any agent derives from the consumption of a unit of good $x$ is precisely $\sqrt{2}$. For the purposes of exchanging the good, agent $A$ would like to utilize a “Dedekind cutter” to mint a medium of exchange such that the numerical value of the medium is precisely identical to that of the utility derivable from the consumption of the good. Although the process by which coins are minted is initially costless, the cost becomes prohibitively high the moment the agent endeavors to mint a sum of coins whose collective face value, when squared, exceeds $2$; moreover, each coin can only express positive rational numbers. Suppose that we have another agent, $B$, who is in possession of the good. Evidently, both agents are willing to perform an exchange, yet they are unable to do so, because one cannot express an irrational number in terms of a finite sum of rational ones. We now consider another mathematically bizarre exchange economy consisting of two agents and two types of goods where the utility anyone can derive from some quantity of good $x$ is a rational multiple of $\sqrt{2}$ and that derivable from some quantity of $y$ is a rational multiple of $\sqrt{\pi}$. Suppose $A$ were in possession of all $x$ and $B$ were in possession of all $y$ and both agents are equally unwilling to allow the other to get more than what is exchanged in return. If $B$ tried to obtain one unit of $x$, she would have to divide her good to an infinite extent, namely, $\sqrt{\frac{2}{\pi}}$, in order for an exchange to occur. It would seem that such exchange economies would paralyze in the absence of “Dedekind cutters”, which themselves are discretely exchangeable; without a sufficiently rigorous description of how bargaining occurs, optimal allocations, despite being guaranteed to exist, are not guaranteed to be reached. Finally, consider an economy where two agents are present and where the first is in possession of an infinitely divisible good, literally the continuum represented by the set $[0,1]$, and the second has the continuum $[2,3]$; it would seem that resources are scarce because only two continua are present. But if we take one real number from $[0,1]$ and give it to the second agent, both agents are indifferent. Agent 1 is not made worse off since $\mathfrak{c}-1=\mathfrak{c}$. Likewise, $\mathfrak{c}+1=\mathfrak{c}$, so the distribution is not Pareto optimal, and permanently so, since we can always make one agent better off without making another worse off.
With this said, it would be natural ask whether it is even possible to axiomatize the foundations of economic theory in terms of logic or mathematics. Should an economic axiom be intrinsically accepted on the basis that it is an empirical observation, such as the assumption of self-interest, diminishing marginal returns or monotonic preferences, despite the fact that it produces unobservable or unfalsifiable conclusions? Or should it be extrinsically accepted because it is capable of accounting for some definite, observable phenomena, despite being intrinsically false? At an academic conference on the discovery of forcing, Paul Cohen noted that
@@
The attempts to formalize mathematics and make precise what the axioms are were never thought of as attempts to explain the rules of logic, but rather, to write down these rules and axioms which appeared to correspond to what contemporary mathematicians were using. [3, p. 1074]
We adopt a similar position with regards to this endeavor to axiomatize a narrow scion of the foundations of economic theory. We certainly do not claim that we are uncovering fundamental, irrefutable truths about economic behavior; we endeavor only to capture and make precise a particular model of the axioms (or hypotheses) that, if true, would produce a set of interesting and non-trivial results; the Coase theorem is one such result.
@@ 3. Foundations
Let $\mathbb{A},R\neq\varnothing$ with $|\mathbb{A}|=n,\;|R|=m$ and $n,m\in\mathbb{N}$. The distribution of resources across each instance of trade $t\in\mathbb{N}$ can be characterized as a mapping $\mathcal{I}_{t}:\mathbb{A}\to\mathcal{P}(R)$ such that the following two conditions are satisfied:
@@
The preferences of each agent can be represented as a mapping $W:\mathbb{A}\to\mathcal{P}\big{(}\mathcal{P}(R)\times\mathcal{P}(R)\big{)}$ such that they are characteristically rational and therefore satisfy the following conditions:
@@
These constructions also preclude transaction costs because no element is lost through the friction of trade.
Lastly, we construct a definition of Pareto optimality.
@@
Given an agent $a_{q}$ and the distributions $\mathfrak{I}=\big{(}\mathcal{I}_{t}(a_{i})\big{)}_{i=0}^{n-1}$, $\mathfrak{K}=\big{(}\mathcal{I^{\prime}}_{t}(a_{i})\big{)}_{i=0}^{n-1}$, we say that the agent strictly prefers $\mathfrak{K}$ over $\mathfrak{I}$ (i.e., ^$\hat{\mathfrak{I}}(a_{q})\prec\hat{\mathfrak{K}}(a_{q})$) iff $\big{(}\mathcal{I}_{t}(a_{q}),\mathcal{I^{\prime}}_{t}(a_{q})\big{)}\in W(a_{q}).$

### v6

@@ 1. Introduction
An externality can be defined as an effect, often negative, that is imposed on a third party through a private activity. An instance of this problem occurs when non-smokers and smokers are in each other’s vicinity; the non-smoker is adversely affected by the smoker’s consumption of cigarettes through no fault of their own. The economic and philosophical significance of the existence of externalities lies in the challenge it poses to the classical position that a free-market economy is capable of allocating resources in a socially efficient manner; if the social costs of private consumption are not reflected in the market price of the product in question, then consumers would be incentivized to consume the good at a level that is beyond socially optimal. A comprehensive overview of the economic explanations surrounding externalities and their historical origins can be found in [2].
Prior to Ronald Coase’s work, The Problem of Social Cost, the conventional view—which Coase himself attributes to Arthur Pigou [5, p. 1]—held that externalities could be corrected through state intervention: namely, through taxes or subsidies that are designed to deter people from either over-consuming goods associated with negative externalities or under-consuming goods associated with positive externalities [12]. Coase reoriented the debate from one that relied solely on governmental regulation to one that recognized the potential efficiency of private negotiation in resolving externalities. In particular, Coase reasoned that externalities resulted from ill-defined property rights; provided that property rights are well-defined and each party could negotiate with the other without hindrance, resources would be distributed away from those who value them less and toward those who value them more irrespective of how the rights are initially distributed [5, p. 8, para. 2]. This proposition, although informal, was characterized as a theorem in George Stigler’s Theory of Price:
@@
Stigler’s recognition of the potential mathematical standing of what is now widely termed the Coase theorem invites us to examine the economic proposition precisely, even though an exact formulation is not present in Coase’s original work. The prevailing interpretation of the theorem adopted by this paper can be stated as follows: if property rights are well-defined and there are no transaction costs, then rational agents would arrive at an optimal distribution of resources of their own accord. We find this interpretation in Stigler [14], Nutter [11], Demsetz [8], and Coase himself [5]. Yet, again, we are confronted with the difficulty of constructing a precise interpretation of informal notions such as the property of being “well-defined” or constituting a “transaction cost”. Additionally, we may interpret the notion of “arriving at an optimal outcome” as a declaration of the existence of some ultimate, optimal distribution such that it is a necessary outcome of every possible initial distribution; alternatively, we may consider that there are multiple ultimate distributions for each initial distribution that fall within the set of outcomes characterized as being optimal.
There have been numerous attempts to formalize the Coase theorem in the literature, and it is essential to understand how the approach in this paper significantly differs from those earlier efforts. In cooperative game-theoretic terms, efficient bargains can be analyzed using the concept of the core of an economy with externalities. In fact, Aivazian and Callen (1981) showed that when more than two parties are involved, the bargaining game may have an empty core, meaning there is no stable allocation that cannot be improved upon by some coalition [1]. In their example, any agreement among a subset of the parties is unstable: other parties can always offer a different deal, leading to what they describe as an “endless recontracting” process. This result implies that with three or more affected agents, private bargaining can fail to produce a stable efficient outcome – a direct challenge to a broad interpretation of the Coase theorem. Coase himself responded to such arguments by suggesting that these theoretical instabilities might not undermine the practical insight of his theorem; the study of such examples, he argued, had “not led [him] to modify [his] views” [6]. Nevertheless, the empty core problem highlights the need to explicitly account for coalition formation and bargaining rules when formalizing Coase’s result.
Other formalizations have focused on the conditions required for Coasean bargaining to yield a unique outcome. Hurwicz [10], for example, proved that quasi-linear utility (i.e. the absence of income effects) is both a sufficient and necessary condition for the final allocation of resources to be invariant to the initial assignment of property rights. In other words, if agents have quasi-linear preferences, then any zero-transaction-cost bargaining leads to the same efficient outcome regardless of who starts with the rights – but if preferences are not quasi-linear, so that wealth effects matter, different initial allocations can lead to different Pareto-efficient outcomes. This finding, sometimes termed Coase’s “neutrality” or invariance theorem, shows that the Coasean prediction of unique outcomes rests on restrictive assumptions. Likewise, formal negotiation models in economics, such as Rubinstein’s alternating-offer bargaining model [13], confirm that two parties with no transaction costs will reach an efficient agreement in equilibrium, but they do not automatically generalize to multiple parties without additional coordination assumptions.
Despite these prior contributions, there remains value in a more explicit logical formalization of the Coase theorem that lays bare the role of each assumption. The approach in this paper differs from earlier work by axiomatizing the bargaining process itself in a set-theoretic framework. Whereas many treatments either assume an equilibrium concept or rely on specific functional forms for utility, here we start from first principles: a finite set of agents, a finite set of resources, and a dynamic rule for voluntary exchanges. This framework allows us to pinpoint precisely why certain interpretations of the Coase theorem can fail. In particular, our model will illustrate, in line with the core literature, that if bargaining is restricted to bilateral trades, the outcome need not be efficient or unique. We then introduce the notion of ideal exchanges – effectively allowing multi-party agreements that improve at least one agent’s lot without hurting others – and show that under this assumption a strong version of the Coase theorem does hold. The additional contribution of this paper is thus a careful analysis of how the bargaining procedure itself affects the validity of Coase’s result, and a demonstration that allowing sufficiently broad cooperation among agents can rescue the theorem.
The remainder of the paper is organized as follows. In Section 2, we discuss the philosophical and methodological motivations for our approach by examining how the Coase theorem is conventionally illustrated and what hidden assumptions these methods entail. This sets the stage for our formal development. In Section 3, we present the formal foundations of our model: the definitions of agents, resources, preferences, and the rules governing exchanges. Section 4 derives some secondary results, showing that two commonly stated versions of the Coase theorem do not hold under the simplest bargaining protocol. Section 5 introduces the concept of ideal exchanges and proves a formal version of the Coase theorem under that assumption. Finally, Section 6 offers philosophical and concluding remarks, reflecting on how these formal results relate to the original spirit of Coase’s insight and the importance of mathematical philosophy in economic theory.
@@ 2. Preliminaries
Before moving to the formal model, it is helpful to review how economists typically illustrate the Coase theorem and to highlight why a more foundational approach might be necessary. Contemporary expositions of the Coase theorem are often informal or rely on geometric arguments that, while intuitively appealing, conceal underlying mathematical assumptions. One common illustration uses the tools of indifference curve analysis in an Edgeworth box, as found in many microeconomics textbooks. We summarize that approach here to ground our discussion.
Consider the existence of two types of goods, $x$ and $y$, such that $Q_{x}$ denotes the quantity of the former and $Q_{y}$ denotes the quantity of the latter. It is standard to assume that each 2-tuple $(Q_{x},Q_{y})$ is present in $\mathbb{R}^{+}\times\mathbb{R}^{+}$ or, equivalently, $\mathbb{R}^{2}_{+}$. [^1] We include 0 in $\mathbb{R}^{+}$. We proceed by declaring the existence of a utility function for some agent $A$ such that $\mathcal{U}_{A}:\mathbb{R}^{2}_{+}\to\mathbb{R}^{+}$ to model the notion that it is possible to map pairs of quantities of goods to some level of utility. [^2] It would seem that happiness is a crude synonym for utility, since the latter is often portrayed as a quantifiable notion. An agent is indifferent between two outcomes iff they render the same level of utility. In other words, an indifference curve is a set of all the 2-tuples $(Q_{x},Q_{y})\in\mathbb{R}^{2}_{+}$ such that $\mathcal{U}_{A}(Q_{x},Q_{y})=c$, where $c$ is a constant in $\mathbb{R}^{+}$. Furthermore, it is standard for an economist to assume that each indifference curve carries important characteristics. Intuitively speaking, it would be peculiar for an indifference curve to contain $(Q_{x},Q_{y})$ and $(Q_{x}^{\prime},Q_{y}^{\prime})$ if $Q_{x}<Q_{x}^{\prime}$ and $Q_{y}<Q_{y}^{\prime}$, since that would imply that the utility of the agent remains constant despite an increase in the presence of both goods. It follows, therefore, that all indifference curves must be monotonically decreasing. A more refined assumption would be that indifference curves should model the empirical observation of diminishing marginal utility, that is, every additional unit of utility gained from each additional unit of the good decreases, but never goes below zero. [^3] If it did, then that would imply that an additional unit of the good resulted in a loss of utility, which is not a standard assumption in microeconomic theory. Thus, indifference curves are often portrayed as being convex to the agent’s origin.
@@
Every Pareto optimal distribution occurs when the indifference curves for both agents are tangent to each other. We can see this by considering the intersection at $\omega_{A}$ or, equivalently, $\omega_{B}$. Since $A$ has an excess of good $x$ and $B$ has an excess of good $y$, it would be possible for $A$ to give up just enough of $x$ and receive just enough of $y$ so that she is indifferent between $\omega_{A}$ and $\Omega$, while $B$’s indifference curve progresses to a point where it is no longer possible to gain any further utility without damaging $A$’s material interests; the opposite is true at $\omega_{A}^{\prime}$. If we assume that exchanges can occur iff they are mutually beneficial, then the negotiated outcome is strictly inside the lens formed by the two indifference curves. Through continual negotiations, the size of the lens shrinks until it is no longer possible to make a mutually beneficial trade. [^4] This is a standard description of how negotiations occur. It is notably present in Buchanan and Tullock’s foundational work on public choice theory, The Calculus of Consent: Logical Foundations of Constitutional Democracy [3, p. 100]. Hence, the ultimate outcome, although different from $\Omega$, would still be Pareto optimal.
@@
It is striking that the above account – standard in microeconomic theory – relies on several mathematical constructs that are taken for granted by economists but raise foundational questions. Edgeworth’s original 1881 formulation of utility in Mathematical Psychics (MP) introduced real-number utility assignments and continuous indifference curves at a time when the rigorous foundations of the real numbers were still being established [9, p. 28]. The broad purpose of MP was to apply mathematical techniques to the moral sciences; in doing so, Edgeworth introduced utility measurements by assigning real numbers to represent individuals’ levels of satisfaction or happiness, facilitating the comparison and aggregation of utility across different people. Furthermore, MP readily and implicitly utilizes the concept of a continuum—suggesting infinitely divisible goods or utility levels—without being concerned with foundational aspects of real analysis. On a historical note, Dedekind’s Stetigkeit und irrationale Zahlen was published in 1872, merely 9 years before MP [7]; it is therefore likely that Edgeworth was unaware of these developments in the philosophical foundations of mathematics and their effects on the mathematical foundations of economics.
[…]
A final thought experiment: suppose two agents exist in a world where one agent’s endowment is literally the continuum of real numbers in the interval $[0,1]$ and the other’s endowment is the continuum $[2,3]$. At first glance, resources seem scarce (each has a finite interval of the real line). But note that these are infinite (uncountable) sets. Agent 1 could give one specific real number (say $0.5$) to Agent 2. After this “trade,” Agent 1 still has $0,1]\setminus\{0.5\}$, which is essentially the interval $[0,1]$ with a single point removed – still uncountably infinite, and in fact equinumerous with the original $[0,1]$ (removing one point from a continuum leaves its cardinality $\mathfrak{c}$ unchanged, since $\mathfrak{c}-1=\mathfrak{c}$). Likewise, Agent 2’s holdings have increased to $2,3]\cup\{0.5\}$, which has the same cardinality as $[2,3]$ ($\mathfrak{c}+1=\mathfrak{c}$). From each agent’s perspective, nothing has materially changed – each still effectively has a “continuum” of resources – so both agents would be indifferent to this trade. By transferring finitely or countably many singletons from $[0,1]$ to the other agent, after each step the first agent still holds a set of cardinality $\mathfrak{c}$, and, under Lebesgue measure, still total measure $1$. Thus, with cardinality as a welfare proxy, these moves never register a loss for the first agent—even though resources are being transferred. However, no countable sequence of singleton trades can exhaust $[0,1]$; only a transfinite process of length $\mathfrak{c}$ could in principle remove all points. This highlights that cardinality is too coarse a notion for welfare.
The bizarre scenario violates the spirit of Pareto optimality – we could keep making one party better off while never hurting the other – yet it arises from taking divisibility and infinite sets seriously without additional structure on preferences or the trading process.
The above examples, while contrived, underscore a deeper point: without a sufficiently careful description of how bargaining occurs, the usual informal arguments for the Coase theorem can run into conceptual puzzles or indeterminacies. If agents can trade in arbitrary real quantities and have exotic preferences, what guarantees that they will reach an efficient outcome? Standard economic reasoning assumes away these difficulties by imposing nice properties such as continuity, convexity, or differentiability, and by implicitly presuming that agents negotiate only in mutually beneficial ways. But to truly prove a Coase-like result, one needs to specify the rules of the game.
This motivates an axiomatic approach. We aim to capture the essence of Coasean bargaining in a simple, logical framework, avoiding unnecessary mathematical structures. Our goal is not to reflect every nuance of real bargaining, but to model a world in which rational agents trade resources if and only if it benefits them to do so, and to see what must be assumed for the Coase theorem to hold in such a world. As Paul Cohen remarked about formalizing mathematics, attempts to write down axioms are often meant to codify the rules practitioners implicitly use:
@@
The attempts to formalize mathematics and make precise what the axioms are were never thought of as attempts to explain the rules of logic, but rather, to write down these rules and axioms which appeared to correspond to what contemporary mathematicians were using. [4, p. 1074]
Here, we attempt something similar for a corner of economic theory: we write down rules that mirror idealized economic assumptions and then explore their consequences. We certainly do not claim these axioms are empirically “true” or that they capture all aspects of real negotiations. Rather, this is a theoretical exercise to determine what follows from particular idealized assumptions about trading behavior. If the Coase theorem is to be regarded as a theorem in the mathematical sense, we must be clear about the assumptions under which it holds and the precise meaning of its conclusion.
@@ 3. Foundations
We now set up a formal model of a bargaining economy to analyze the claims of the Coase theorem. The model consists of a finite set of agents, a finite set of divisible resources, and a discrete time process during which agents may exchange resources. The rules of exchange will be explicitly defined to capture the idea of voluntary, mutually beneficial trades in a world without transaction costs. This framework is deliberately simplified, consisting of finite sets, no production or uncertainty, so that we can focus on the logical structure of Coasean bargaining.
Let $\mathbb{A}$ be a nonempty finite set of agents (with $|\mathbb{A}|=n$ agents) and let $R$ be a nonempty finite set of resources (with $|R|=m$ distinct divisible goods or items). For convenience, we label the agents as $a_{0},a_{1},\ldots,a_{n-1}$, and we will often refer to an agent $a_{i}\in\mathbb{A}$ by index $i$. A distribution of resources at any given time $t$ is a specification of which agent holds each resource. Formally, we represent a distribution as a function $\mathcal{I}_{t}:\mathbb{A}\to\mathcal{P}(R)$, where $\mathcal{P}(R)$ denotes the power set of $R$ (the set of all subsets of $R$). We impose two natural conditions on any distribution $\mathcal{I}_{t}$:
@@
Each agent has preferences over possible bundles of resources. We represent the strict preference relation of agent $a_{i}$ by a set $W(a_{i})$ consisting of ordered pairs of bundles. If $(A,B)\in W(a_{i})$, it means that agent $a_{i}$ strictly prefers bundle $B$ over bundle $A$. Formally, let $W:\mathbb{A}\to\mathcal{P}(\mathcal{P}(R)\times\mathcal{P}(R))$ be a function assigning to each agent $a_{i}$ a set $W(a_{i})$ of ordered pairs $(A,B)$ with $A,B\subseteq R$. We assume each $W(a_{i})$ satisfies the standard rationality properties of a strict total order:
@@
In simple terms, each agent’s preference $W(a_{i})$ is a strict total order over the set of all possible bundles of resources. We can think of each $W(a_{i})$ as generated by a utility function $u_{i}:\mathcal{P}(R)\to\mathbb{R}$ that assigns a numerical utility to each bundle – then $(A,B)\in W(a_{i})$ if and only if $u_{i}(A)<u_{i}(B)$. However, we do not require the existence of a numeric utility representation, as the axioms (B1)(B3) are just the ordinal preference assumptions that ensure a well-defined ranking.
@@
In simpler terms, (C1) says “if two agents have what each other wants and both would be happier after swapping, then they will swap,” and (C2) says “if no such swap exists, nothing changes.” These axioms formalize the idea of frictionless bargaining where any opportunity for a Pareto-improving bilateral trade is immediately seized. They also implicitly assume no transaction costs: exchanges do not consume any resources or utility beyond the goods being swapped, and no goods vanish in the process. Thus, if a beneficial trade is possible, it happens and costs nothing; if this is not possible, the status quo persists. We emphasize that conditions (C1) and (C2) are inductive rules governing the evolution of allocations over time. Starting from $\mathfrak{I}_{0}$, (C1) and (C2) determine $\mathfrak{I}_{1}$, which in turn provides the basis to find trades for $\mathfrak{I}_{2}$, and so on. Because the set of possible distributions is finite, as there are only finitely many ways to assign $m$ goods to $n$ people, this process cannot continue indefinitely without repeating a prior distribution. We will shortly prove that the process indeed converges to a stable allocation in a finite number of steps.
Lastly, we define more carefully what it means for a distribution of resources to be Pareto optimal. Intuitively, an allocation is Pareto optimal if there is no alternative allocation that would make some agent better off without making any other agent worse off. In our formal setting, this means: a distribution $\mathfrak{I}$ is Pareto optimal if whenever there is at least one agent who strictly prefers some other distribution $\mathfrak{K}$ to $\mathfrak{I}$, then there must be at least one different agent who strictly prefers $\mathfrak{I}$ to $\mathfrak{K}$. Thus, you cannot move to $\mathfrak{K}$ without hurting someone, if $\mathfrak{I}$ is Pareto optimal.
@@
Given an agent $a_{q}$ and the distributions $\mathfrak{I}=\big(\mathcal{I}_{t}(a_{i})\big)_{i=0}^{n-1}$, $\mathfrak{K}=\big(\mathcal{I^{\prime}}_{t}(a_{i})\big)_{i=0}^{n-1}$, we say that the agent strictly prefers $\mathfrak{K}$ over $\mathfrak{I}$ (i.e., $\hat{\mathfrak{I}}(a_{q})\prec\hat{\mathfrak{K}}(a_{q})$) iff $\big(\mathcal{I}_{t}(a_{q}),\mathcal{I^{\prime}}_{t}(a_{q})\big)\in W(a_{q}).$